OpenStudy (anonymous):

instructions: Solve each equation for P. MUST SHOW WORK!!! PLEASE READ INSTRUCTIONS!!! 1. xp+yp=z 2. n=p-k/j 3. a=b+cp 4. p+3/m = -1 instructions: solve each equation for the given variable. 1. -2z - xy = x + 7 for x 2. a/b - 8 = c/d for a

OpenStudy (allank):

Why does this look like a homework question? :)

OpenStudy (lgbasallote):

it is

OpenStudy (anonymous):

yah

OpenStudy (allank):

@Igbasallote :)

OpenStudy (anonymous):

i'm having trouble understanding it, my teacher didn't have enoough time to explain in class and this is only a workbook

OpenStudy (anonymous):

p(x)+p(y)=z p(x+y)=z p(x+y)/x+y=z/x+y p=z/x+y

OpenStudy (anonymous):

well now you make it look simple, pisses me off smart people lol THANKS FOR THE HELP!!! are the rest of the questions able to solve this way?

OpenStudy (anonymous):

i mean the second portion?

OpenStudy (anonymous):

Yeah, for example for the first one basically just put all the x variables in one side factor out y and simplify.

OpenStudy (anonymous):

thanks for the help, god i hate algebra!